Is this a trick question? I am not sure how to do these. Can some show me what steps I am suppose to take for this.
Solve each equation:
a. X2 = -100
b. 5X2 = -125

Answer by stanbon(75887)

a. X^2 = -100
There are no Real Numbers whose square is negative.
There are Complex Numbers that do the job.
For your problem
X= 10i or -10i where "i" is the square root of -1.
----------------------
b. 5X^2 = -125
X^2=-25
X=5i or X=-5i
---------------
Cheers
Stan H.